What is the area, in square feet, of a rectangle whose width is 6 feet and whose length is 2 feet more than three times it's width?
Answer by macston(5194)

width=W=6 ft; length =L=3W+2=3(6)+2=20 ft
Area=LW=(6 ft)(20 ft)=120 sq ft